# A Comprehensive Survey on Network Traffic Synthesis: From Statistical Models to Deep Learning

### v1

@@ Abstract
Synthetic network traffic generation has emerged as a promising alternative for various data-driven applications in the networking domain. It enables the creation of synthetic data that preserves real-world characteristics while addressing key challenges such as data scarcity, privacy concerns, and purity constraints associated with real data. In this survey, we provide a comprehensive review of synthetic network traffic generation approaches, covering essential aspects such as data types, generation models, and evaluation methods. With the rapid advancements in AI and machine learning, we focus particularly on deep learning-based techniques while also providing a detailed discussion of statistical methods and their extensions, including commercially available tools. Furthermore, we highlight open challenges in this domain and discuss potential future directions for further research and development. This survey serves as a foundational resource for researchers and practitioners, offering a structured analysis of existing methods, challenges, and opportunities in synthetic network traffic generation.
@@ I Introduction
With the unprecedented growth of network traffic and the variety of network-based applications in recent years, traffic models and traffic traces have become essential components for both network providers and users for tasks such as network design, capacity planning, Quality of Service (QoS) provisioning, performance tuning, and optimization. Accurate and representative traffic plays a significant role in understanding network behavior [165, 74], maintaining QoS [91, 15, 96], optimizing systems [126, 157], and benchmarking critical network systems [115]. Furthermore, network traffic analysis is crucial for securing network systems while protecting user privacy [67, 154, 131], identifying bottlenecks in real-time [6, 24], and forecasting future demands [4, 41, 114] to name a few.
@@
While measured traffic traces and those generated using statistical models served as a main pillar of network traffic analysis in the past, the emerging design, provisioning, and management tools are based on data-driven approaches such as machine learning (ML) and artificial intelligence (AI). A major challenge for developing and training such modes is the need for large and complex data sets that represent network traffic including the interaction of myriads of technologies, protocols, and applications. For example, Deep Neural Network (DNN)-based tools typically require large volumes of network traffic data to achieve optimal performance [136, 67]. However, it is essential that network traffic data, rich in fundamental traffic characteristics, provide valuable insights into their behavior. Another example where synthetic traffic will play a critical role is Digital Twin technology, which is envisioned to be an enabler for innovation in 5G/6G networking in many scenarios including industrial IoT and V2X systems  [101, 84]. In many circumstances, creating the necessary conditions on physical networks for gathering appropriate data for such systems may not even be possible. For example, a digital twin may be developed prior to the actual deployment of a novel 6G configuration, or one may need to create multiple disruptions to evaluate the robustness of a configuration. Such cases can be handled only via synthetic traffic traces. Even with a dedicated network, collecting high-quality traffic from real-world settings remains challenging.
@@
First, the sheer volume of data transmitted through networks makes it difficult to extract relevant traffic while filtering out noisy background data. Second, deploying data collection tools in complex network environments often requires extended periods, leading to delays in completing network traffic analysis tasks [36]. Third, controlling the traffic sources to operate a network under the necessary conditions for trace gathering or for creating anomalies may interfere with normal operation or even be impossible to achieve. Additionally, stringent data regulations and privacy policies further complicate the process. These regulations often restrict operators and researchers from freely collecting, sharing, or analyzing traffic data, as doing so could expose sensitive information [119, 54].
To address these challenges, researchers have increasingly turned to synthetic data generation. The use of synthetic data to meet data challenges is not unique to networking, as can be seen in its success in fields such as medical image-based diagnosis [180, 29], object detection on autonomous vehicles [72] and fault detection [160, 20]. This approach has been proven instrumental in achieving optimal and robust outcomes in network traffic analysis. Synthetic network traffic generation frameworks are designed to create artificial datasets that preserve the characteristics of real-world data while mitigating issues such as data privacy, security concerns, and noise in the data. By leveraging synthetic data, either independently or in combination with limited real-world data, many network traffic analysis tasks (such as aforementioned) have demonstrated enhanced performance [137, 136, 68].
@@
Reviewing the recent research literature, we observe that a plethora of work has been done in this domain, which we identify under two primary categories. The first category includes early methods based on statistical models, which analyze the statistical characteristics of network traffic features and generate new data that conform to these properties. The second category consists of generative AI models, such as Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), Diffusion Models (DMs), and transformer-based techniques. These methods, supported by deep neural networks (DNNs), have significantly advanced synthetic traffic generation by capturing complex network traffic patterns from training data and replicating them in synthetic traces.
Additionally, we identify and review work that extends these two approaches to network traffic simulators, which often integrate statistical or ML/AI tools to generate network packet streams while preserving protocol-specific attributes. Furthermore, some work has been extended to commercially available network traffic generators that incorporate configurable parameters and are applied in both research and industrial settings.
To highlight the growing interest in the field of network traffic generation, Fig. 1 presents the number of publications in this domain from 2000 to 2024, categorized into four key approaches: statistical methods, Deep Learning (DL) techniques, network simulation, and commercially available generators. The articles were retrieved through a Scopus search using relevant keywords in the titles, abstracts, and keyword sections. Notably, since 2018, there has been a marked shift toward DL-based methods, as evidenced by a sharp increase in publication rates over the past five years.
@@
Figure 1: Research papers in network traffic generation
Motivated by these findings and the growing interest in the field, this survey paper provides a detailed examination of advancements in network traffic generation. To our knowledge, it is the first survey to systematically explore and compare a wide range of synthetic network traffic generation techniques, including traditional statistical methods, DL-based models, network traffic simulators, and commercially available traffic generators. This analysis provides a broader view compared to earlier surveys, which mainly focused on specific generative models like GANs [3, 118]. For example, we begin by reviewing the statistical approaches, followed by Machine Learning-based models, like Autoregression and K-nearest neighbor clustering. Next, we examine DL-based methods, giving more attention to this area due to the growing interest and their ability to handle complex traffic patterns including those for time-varying networks. These include Autoencoders, GANs, DMs, and Transformer-based architectures. We also look at different data formats and several open challenges in network traffic generation followed by future directions.
To summarize, our contributions are as follows:
  • Comprehensive review across different methods: We provide a detailed analysis of various network traffic generation approaches, spanning from statistical methods and DL-based models to extensions into network traffic simulators and commercially available network traffic generators. Through this, we provide a broad picture of how these different methods have evolved in terms of model architectures being used in data generation.
  • Analysis of different data types used: We provide a detailed taxonomy of different data types of interest for network traffic generation. This analysis is particularly based on what information resides in data samples and how the real data is collected at different network layers.
  • Identification of open challenges: We also identify several open challenges, particularly regarding aspects such as privacy concerns, computational efficiency and complex protocol standards, etc. This discussion provides guidelines for further growth in research in this field.
  • Exploration of future research directions: We discuss the potential commercialization of advanced generative models, along with the need for synthetic data techniques to handle unseen network scenarios for enhanced system resilience.
The remainder of the paper is organized as follows: existing surveys related to this research area are reviewed in Section II. Section III discusses various data types used in network traffic generation. The review continues with an exploration of statistical methods in Section IV, followed by an in-depth analysis of deep learning-based approaches in Section V. We present simulations and commercially available tools in Section VI. Open challenges in network traffic generation are addressed in Section VII. Finally, in Section VIII, we provide hints for future research directions, and concluding remarks are provided in Section IX.
II Related Surveys
Previous surveys have largely centered around the use of GANs for generating network traffic data across a wide range of applications, with GANs long being considered state-of-the-art (SOTA) for such tasks. For example, the authors in [23] reviewed how GANs have been used for network traffic generation, categorizing network traffic under time-series patterns and distinguishing it from other types such as flow and sequence generation. This work highlights GANs’ effectiveness in simulating network behaviors with time-dependent characteristics. Bourou et al. [21] reviewed the use of GANs for tabular data synthesis, particularly for generating intrusion detection system (IDS) datasets, where GANs were shown to model the behavior of network attacks effectively. Similarly,  [9] explored the evolution of GANs, demonstrating their role in generating network flows and showcasing them as critical tools for future developments in both network traffic generation and the broader field of cybersecurity systems.
Building on these insights,  [3] assessed the application of GANs in creating synthetic attack data, emphasizing their contribution to building more robust cybersecurity defenses. This line of work positions GANs as essential for generating high-quality data that can be used for stress-testing security systems. The authors in  [56] took a broader approach by conducting a systematic review of generative ML models, including GANs, autoencoders, and hybrid models, and investigated how these models could be leveraged for intrusion detection. Moreover, the work in [118] conducted a comparative study focusing on different GAN architectures, exploring their effectiveness in generating malicious cyber-attack data. The study identified the strengths and limitations of different GAN variants in replicating diverse attack behaviors.
In parallel,  [2] provided a comprehensive survey of network traffic generators, categorizing them based on their functional behaviors and classifying how researchers have used these tools across different network simulation scenarios. This broader focus, while still heavily centered on GANs, also touched upon other data generators, highlighting their growing role in network traffic analysis and performance testing.
@@ II Related Surveys
In contrast to existing surveys on network traffic generation, which have primarily focused on specific model types, particularly the GANs, this review broadens the scope by analyzing a diverse range of generative models, including statistical methods, DL approaches, and simulation-based platforms. In particular, we cover recent advancements with SOTA generative models, such as DMs and Transformers, which have not been addressed in previous surveys. Additionally, it examines critical aspects such as data types, evaluation metrics, and open challenges in the field. Furthermore, we provide insights into how the network traffic generation domain is evolving and advancing toward the future.
III Data types in network traffic generation
@@
Figure 2: Different network traffic data types/formats used and data collection points in network traffic generation are categorized according to the network OSI layer.
@@
Figure 3: Extraction of payload, packet level, flow level, and raw signal using a .pcap. The sample .pcap is from [159]
In this section, we provide a detailed overview of different data formats used in network traffic generation. We further study different data collection tools used in network vantage points. Fig. 2 shows the overview of this categorization according to different network OSI layer and taking a sample packet structure to indicate different information extract at each layer.
III-A Different data types or formats
We identify three main data formats in network data generation namely: Payload, Packet header data, Raw signal and Network usage conditions.
III-A1 Payload
In payload generation, actual byte values related to the message carried in the network packet (i.e., added at the Application layer) is generated. These byte values are typically represented as a word (2 bytes) in hexadecimal notation on tools such as Wireshark [159], Tcpdump [151] as shown in Fig. 3. In literature we observe that generation of payload has got less attention compared to other data types (e.g., packet header data) due to two main reasons. First, byte value patterns in payload data is highly complex with higher variance. Therefore, training a generative AI model that can learn these complex patterns is significantly challenging. Even if such patterns are generated, they cannot be used for reconstructing original application data. Second, for many network related downstream tasks, having packet header information is sufficient to get accurate analytics outcomes. For example, any changes in the payload are replicated in packet length parameter in the packet header. Note that we have not found major related work that cover HTTP/HTTPS data generation that mainly comes under Application layer protocols.
Despite the aforementioned challenges, the paper, PACGAN [30] proposes a GAN-based payload generation approach. In this approach, authors first separate the two digits of the hexadecimal notation of a byte. Then, each digit is mapped to a sub-range of serial values (i.e., one-to-many mapping) which will be fed to the GAN model for as the input. The output sequences are re-mapped to the original hexadecimal notations to create the byte values. Each generated byte value is then duplicated to $d\times d$ blocks forming a large matrix by concatenating each block. The final byte values are derived taking the mean of the blocks.
In [55], Guo et al. pre-process the data by removing protocol headers and extracting the first 784 bytes from the application layer. Packets exceeding 784 bytes are truncated, while those smaller are padded with 0x00 values. The authors argue that focusing on the payload retains the packet’s original information, and limiting it to the first 784 bytes ensures real-time performance. A similar method to generate payload is proposed in [156], however, the authors have fixed the length of the byte vector to 1480 bytes.
We also observe that the simulation based network traffic generators including NS3 [7], iperf [129], ositano [121], TRex [33], and Scapy [147], also create data packets with payloads. While some tools, such as Scapy, enable users to customize payloads by entering specific messages, others generate payloads randomly, simply following the rules of different network protocol. This is because, the main purpose of these simulators is to test the network performance, for example, to measure key metrics such as bandwidth, latency, and jitter and test the speed and reliability of network links, optimize configurations, etc.
III-A2 Header based generation
This is the most common network packet component used for data generation. Observing literature, we identify three main reasons why header-based generation is important compared to other data formats (e.g., payload, bit sequences from physical layers, etc.). First, packet header contains summary statistics of the network packet. For example, packet length or packet size, protocol information, and packet arrival time are some of those key information that we can easily extract from the header. Second, this header information is useful and sufficient for many downstream tasks, particularly in ML-driven analysis of network traffic data. For example, ML-based network traffic classification models to identify the usage of different applications [94, 64], user activities [39] etc., take these packet header information as the input. Third, with the availability of range public (e.g., Wireshark,  [159], Tcpdump [151]) and proprietary tools (e.g., SolarWinds Network Performance Monitor [141], Cisco Meraki  [34]), extraction and monitoring of unencrypted packet header information is easy.
Information extracted at different layers:  As shown in Fig. 2, there are three main network OSI layers used for packet header extraction and used for network traffic generation. While the packet header from each layer contains various parameters, we highlight only the mostly used key information for data generation below.
  1. 1.
    Transport layer: Commonly known as TCP/UDP header or datagram header, Transport layer packet header contains source port, destination port, Flags (SYN, ACK, FIN, and RST) [64, 171].
  2. 2.
    Network layer: Also known as segment header. The total length of the packet, source and destination IP addresses, time-to-live (TTL) and protocol field in the upper layer (e.g., TCP or UDP) are the commonly used Network layer information for data generation [171, 167].
  3. 3.
    Data link layer: Source and destination MAC (Medium Access Control) addresses, ether type (i.e., protocol encapsulated in the upper layer) and frame length are the common features extracted [40].
While several works [137, 136] have focused on generating data extracted from a single layer (e.g., frame length from the Data Link layer), other research has also explored generating data from multiple layers. For example, in  [171, 64], the authors generate multivariate data by extracting source and destination IP addresses and protocol information from the Network layer, along with source and destination port data from the Transport layer.

### v2

# A Comprehensive Survey on Synthetic Network Traffic Generation

@@ Abstract
Synthetic network traffic generation has emerged as a promising alternative for various data-driven applications in the networking domain. It enables the creation of synthetic data that preserves real-world characteristics while addressing key challenges such as data scarcity, privacy concerns, and purity constraints associated with real data. In this survey, we provide a comprehensive review of synthetic network traffic generation approaches, covering essential aspects such as data types and generation models. With the rapid advancements in Artificial Intelligence (AI) and Machine Learning (ML), we focus particularly on deep learning (DL)-based techniques while also providing a detailed discussion of statistical methods and their extensions, including commercially available tools. We present a comprehensive comparision of generation approaches and provide an AI tool to apply this comparision for any network traffic generation papers. Furthermore, we highlight open challenges in this domain and discuss potential future directions for further research and development. This survey serves as a foundational resource for researchers and practitioners, offering a structured analysis of existing methods, challenges, and opportunities in synthetic network traffic generation.
@@ I Introduction
With the unprecedented growth of network traffic and the variety of network-based applications in recent years, traffic models and traffic traces have become essential components for both network providers and users for numerous tasks. For example, accurate and representative traffic plays a significant role in understanding network behavior [188, 88], maintaining Quality of Service (QoS) [106, 15, 112], optimizing systems [145, 177], and benchmarking critical network systems [134]. Furthermore, network traffic analysis is crucial for securing network systems while protecting user privacy [80, 171, 151], identifying bottlenecks in real-time [6, 28], and forecasting future demands [4, 45, 133] to name a few. However, the need to depend on learning-based strategies to tackle such issues under complex scenarios has renewed interest in synthetic traffic generation.
I-A Need for accurate network traffic data and challenges
@@
While measured traffic traces and those generated using statistical models served as a main pillar of network traffic analysis in the past, the emerging design, provisioning, and management tools are based on data-driven approaches such as Machine Learning, Artificial Intelligence and simulated environments such as Digital Twin (DT). A major challenge for developing and training such models is the need for large and complex datasets that represent network traffic including the interaction of myriads of technologies, protocols, topologies, and applications. For example, Deep Neural Network (DNN)–based tools typically require large volumes of network traffic data to achieve optimal performance [156, 80]. Similarly, DT technology—an enabler for simulating 5G/6G networking in industrial Internet of Things (IoT), testing network disruptions and developing Vehicle-to-Everything (V2X) systems [117, 99] etc., relies on accurate network traffic data even before real-world implementations.
Even with dedicated systems and tools, collecting high-quality network traffic from real-world settings remains challenging due to multiple reasons. First, the sheer volume of data transmitted through networks makes it difficult to extract relevant traffic while filtering out noisy background data. In particular, these networks and the related traffic are highly diverse in their characteristics (e.g., protocol, burstiness, packet sizes etc.). For example, wired backbone networks often display high-throughput and relatively stable flows, whereas wireless and cellular networks are influenced by factors such as signal variability, interference, and user mobility, leading to more bursty and dynamic traffic patterns.
@@
Second, deploying data collection tools in complex network environments often requires extended periods, leading to delays in completing network traffic analysis tasks [41]. Third, controlling the traffic sources to operate a network under the necessary conditions for trace gathering or for creating anomalies may interfere with normal operation or even be impossible to achieve. Additionally, stringent data regulations and privacy policies restrict operators and researchers from freely collecting, sharing, or analyzing traffic data [139, 64].
I-B Significance of synthetic network traffic generation
To address these challenges and support diverse network applications, researchers have increasingly turned to synthetic network traffic generation, a trend further motivated by the success of synthetic data in other domains. For instance, synthetic data has been widely adopted in medical image–based diagnosis [207, 32], object detection for autonomous vehicles [86], and fault detection in industrial systems [182, 24].
The primary aim of synthetic network traffic generation frameworks and models is to create artificial datasets that preserve the key characteristics of real-world traffic. These frameworks are typically based on mechanisms such as statistical approaches, deep learning–based methods, or commercially developed traffic simulation tools, each capable of generating diverse types of network traffic data. By leveraging synthetic data, either independently or alongside limited real-world traces, many network traffic analysis tasks have demonstrated enhanced performance [157, 156, 81], highlighting the value of synthetically generated data. For example, it has been already proven that many downstream network traffic analytics tasks including ML-based network traffic classification [157, 110], traffic modeling [11], network security related activities [154, 142], and network simulation development [191, 87] etc. have been benefited through synthetic traffic data. To support these requirements, synthetic network traffic generation acquire positive attributes in various aspects as below.
  • Data fidelity: Modern generative AI models (e.g., DNN-based) can capture subtle variations in network traffic and generate synthetic data that closely mirrors the properties of real-world traffic.
  • Data diversity: These models can produce network traffic for a wide range of network data types, e.g., from local area networks (LAN) to wired backbone networks.
  • Safety and security: Enables controlled emulation of rare failures, attacks, and anomaly conditions for training and evaluating defense mechanisms without disrupting production networks.
  • Benchmarking and reproducibility: Provides standardized, shareable datasets to compare algorithms fairly across research groups and vendors.
  • Pre-deployment validation: Supports stress testing of new architectures (e.g., 6G, IoT, V2X) and closed-loop Artificial Intelligence for IT Operations (AIOps) pipelines before real traffic is available.
  • Privacy and compliance: Facilitates experimentation while avoiding exposure of sensitive user data and easing compliance with data protection regulations.
  • Cost and coverage: Reduces data collection costs and enables exploration of corner cases that are rare or infeasible to capture in the wild.
  • Time efficiency: Compared to deploying hardware or software data collection tools, which require prolonged time, these models can be trained with limited amount of already available data relatively taking less time.
@@ I-B Significance of synthetic network traffic generation
Reviewing the recent research literature, we observe that a plethora of work has been done in this domain, which we identify under two primary categories: i) Statistical models, which analyze the statistical characteristics of traffic features and generate new data that conform to these properties. ii) DL-based models (Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), Diffusion Models (DMs), and transformer-based techniques) supported by DNNs. These models have shown remarkable performance in learning subtle and complex traffic patterns to generate high-fidelity data. Extending the models in these two approaches, network traffic simulators exist that generate network packet streams while preserving protocol-specific attributes. Furthermore, commercial traffic generators, with configurable parameters are applied in both research and industrial settings.
To highlight the growing interest and, hence, the significance in the field of network traffic generation, Fig. 1 presents the number of publications in this domain from 2000 to 2024 under the aforementioned categories. The articles were retrieved through a Scopus search using relevant keywords in the titles, abstracts, and keyword sections. we observe a sustained increase in publications across all categories over the last two decades, with the steepest growth in DL-based methods after 2018. Simulation- and statistical-based approaches remain active but comparatively flatter, indicating methodological maturity and the inadequate scalability and configurability to meet the emerging needs, while the steady rise of commercially used tools since the mid-2010s signals growing industrial adoption. The overall peak in 2023–2024 underscores that the importance of traffic generation is increasingly recognized by both researchers and practitioners.
@@
Figure 1: Research papers in network traffic generation. The numbers were extracted by querying the Scopus database.
I-C Related Surveys
Recently, a growing body of work has explored synthetic data generation across diverse domains, including several surveys that examine common challenges in data-driven generation. Surveys by Goyal et al. [62], Bengesi et al. [19], and Figueira et al. [56] examine the topic from a general ML perspective, addressing cross-cutting challenges such as data scarcity, privacy, and algorithmic bias rather than focusing on specific applications. The surveyed models span traditional ML techniques to advanced DL approaches such as GANs, VAEs, diffusion models, and Generative Pre-trained Transformers (GPTs). Figueira et al. [56] provide a detailed overview of GAN-based data generation for tabular datasets, highlighting key advancements, evaluation strategies, and training challenges.
Table I: Comparison of existing survey papers on synthetic network traffic generation
Survey Methodologies or Models Covered Scope of Survey Limitations Identified
[27] GANs Time-series modeling of network traffic Limited to GANs and time-series traffic patterns
[25] GANs Intrusion Detection System (IDS) dataset generation using tabular data Does not generalize beyond tabular attack simulation
[9] GANs Evolution and application of GANs for network flows No coverage of other generative methods
[3] GANs GAN-based attack simulation for cybersecurity testing GAN-centric, no discussion on evaluation or metrics
[67] GANs, Autoencoders, Hybrid Models Generative models for IDS Broader model types, but focus still limited to IDS context
[138] GAN Variants Comparative study of GANs for cyberattack data Focused only on malicious attack traffic and GANs
[2] Simulation tools, GANs Functional categorization of traffic generators No deep technical breakdown of DL-based generators
Collectively, these studies establish a strong methodological foundation for synthetic data generation. Building on this foundation, numerous works have focused on specific applications such as image generation [120] and time-series generation [72] etc. In the context of network traffic generation, the related surveys (see Table I) have primarily examined methodologies that apply generative AI models to capture complex temporal and behavioral patterns in network environments. We note that the majority of existing surveys on network traffic generation are limited to GAN-based approaches, although they span diverse applications such as cyberattack simulation and generic time-series traffic modeling etc.
I-D Contributions of this survey
Motivated by these findings and the growing interest in the field, this survey paper provides a detailed review of network traffic generation. To our knowledge, it is the first survey to systematically explore and compare a wide range of synthetic network traffic generation techniques.
@@ I-D Contributions of this survey
In contrast to tprevious surveys on network traffic generation, which have primarily focused on specific model types, particularly the GANs, this review broadens the scope by analyzing a diverse range of generative models, including statistical methods, DL approaches, and simulation-based platforms. In particular, we cover recent advancements with state-of-the-art (SOTA) generative models, such as DMs and Transformers, which have not been addressed in previous surveys. We also review different network traffic type data generation to provide broader understanding about the nature of data handling with generative models. In addition to existing literature, we discuss open challenges (e.g., privacy-utility tradeoff, complexity, scalability etc.) in network traffic generation with plausible solutions as future research directions.
In this review, we adopt a network- and application-agnostic approach, where the collected network traffic can originate from diverse environments (e.g., LANs, enterprise systems, IoT networks) and support various applications (e.g., cybersecurity, addressing class imbalance). Since most of these networks follow the standard layered architecture, the corresponding data types and protocols (e.g., payload, packet- and flow-level data, physical-layer signals) are similar, and thus the traffic generation process remains consistent. To summarize, our contributions are as follows:
  • Comprehensive review across different methods: We provide a detailed analysis of various network traffic generation approaches, from statistical methods and DL-based models to extensions into network traffic simulators and commercially available traffic generators. We provide a broad picture of evolvement of the models in terms of their architectures used in data generation.
  • Comparative analysis of major models: We provide a comparison between major DL-based models using a range of metrics, further guiding users for appropriate model selection for their generative tasks.
  • Analysis of different data types used: We provide a detailed taxonomy of different data types of interest for network traffic generation. This analysis is particularly based on what information resides in data samples and how the real data are collected at different network layers.
  • Identification of open challenges: We also identify several open challenges, particularly regarding aspects such as privacy concerns, computational efficiency and complex protocol standards, etc. This discussion provides insights for further growth in research in this field.
  • Exploration of future research directions: Following the open challenges, we provide plausible solutions to overcome those challenges as a guidance to the research community to advance further research.
Figure 2: Overview of the sections
The remainder of the paper is organized as follows: as shown in Fig. 2, we organize existing techniques into a systematic taxonomy with a structured flow between them. In Section II, we explain the different data types primarily used for network traffic generation, which influence the design and development of the underlying mechanisms (e.g., data pre- and post-processing) in the overall generation pipelines.
Building on these data sources, we first introduce statistical approaches for network traffic generation in Section III. After highlighting their limitations, we present the DL-based approaches in Section IV, which in general are more advanced tools for network traffic generation that address limitations of statistical methods. We subsequently analyze the distinctions between statistical and deep learning–based approaches and outline a constraint-driven strategy for model selection in Section V. Leveraging the knowledge from previous sections, we then describe simulation-based methods in Section VI. Commonly employed in commercial settings and incorporate both statistical and DL-based techniques as internal mechanisms. Finally, in Section VII, we discuss open challenges in network traffic synthesis and provide insights for future research directions, and then conclude the paper in Section VIII.
II Data types in network traffic generation
@@
Figure 3: Network traffic data representations and data collection points organized according to the OSI model. Each layer exposes distinct observable data that can be used for traffic generation at packet or flow granularity.
@@
Figure 4: Extraction of packet- and flow-level representations from a .pcap file.
In this section, we characterize the network traffic data used for synthetic traffic generation. Traffic collected from different network vantage points provides different levels of data visibility. For instance, application-layer proxies observe application-level information, often in encrypted form, whereas network routers expose IP addresses, port numbers, and other header fields. At the physical medium, monitoring captures raw physical signals, such as binary sequences or electrical or wireless signals. Consequently, monitoring locations across the network stack expose different subsets of traffic information, leading to substantial variation in the data available for synthetic traffic generation.
To systematically organize these observations, we map each data representation to the Open Systems Interconnection (OSI) model, which clearly characterizes observable network information according to monitoring vantage points. Existing traffic generation models operate on data extracted at different OSI layers and at varying levels of aggregation. Accordingly, we structure the discussion layer by layer, from the Application layer (L7) to the Physical layer (L1). We further distinguish between packet-level and flow-level representations as alternative aggregation views over the same underlying observations. Fig. 3 summarizes the relationship between network vantage points, OSI layers, observable data types, and common data collection mechanisms.
II-A Application layer (L7): Payload and application semantics
At the Application layer, network traffic is represented by the payload, i.e., the sequence of bytes generated by applications, encrypted and encapsulated within packets. In packet capture tools such as Wireshark [181] and Tcpdump [167], these bytes are typically visualized in hexadecimal format (Fig. 4). From a data visibility perspective, full payload access depends strongly on the monitoring vantage point. When traffic is collected at end hosts, proxies, or application gateways, payload contents may be directly observable. In contrast, when monitoring is performed at intermediate network points such as routers, switches, or backbone links, payloads are increasingly encrypted, limiting visibility to encrypted byte streams or metadata only. As this survey focuses on traffic generation from the perspective of these intermediate nodes, the analysis of unencrypted application data is considered out of scope.
Across the literature, application-layer payload data has received comparatively limited attention in traffic generation studies. This is primarily due to two factors. First, modern network traffic is predominantly encrypted (e.g., HTTPS, QUIC), which prevents direct observation of application content at most network vantage points. As a result, realistic payload data can typically only be obtained through end-device instrumentation, trusted middleboxes, or controlled testbed environments. Second, even when available, payload data exhibits high variability and complex semantic structure, making it challenging to model and often unnecessary for many downstream tasks, where header-level features such as packet size, timing, and protocol metadata already provide sufficient discriminative power.
Despite these challenges, a small number of studies have explored payload-based traffic generation by operating directly on application-layer byte sequences [33, 65, 176]. In these works, payloads are typically extracted after removing protocol headers and converted into fixed-length byte vectors through truncation or padding, enabling consistent representation across packets. This allows models to preserve fine-grained application semantics embedded in the payload, while still remaining constrained by practical considerations such as limited payload visibility in encrypted traffic, data availability, and computational overhead.
Data collection mechanisms: Packet capture tools such as Wireshark [181] and Tcpdump [167] can record payload data in raw or hexadecimal format when traffic is unencrypted or captured at trusted endpoints (Fig. 4). Passive monitoring frameworks such as Zeek [168], however, typically do not retain full payloads and instead extract and log application-layer metadata, including Uniform Resource Identifiers (URIs), HTTP headers, and Multipurpose Internet Mail Extensions (MIME) types. Deep Packet Inspection (DPI) systems can collect and analyze payloads directly, most effectively in unencrypted traffic, using techniques such as regular expression matching and grammar-based parsing to identify protocols, extract features, and detect anomalies. In encrypted settings, payload visibility is limited, and only derived metadata or side-channel features are typically available for traffic analysis and generation.
II-B Transport, Network, and Data Link layers (L4–L2): Packet headers